a bowling ball of mass 7kg is thrown with a speed of 11m/s at a single bowling pin of mass 1.5kg. Calculate the speed the pin moves at after the collision, if the ball continued at a speed of 6m/s


1
Expert's answer
2021-09-14T09:35:48-0400

"MV=MV'+mv'\to v'=(MV-MV')/m="


"=(7\cdot 11-7\cdot6)/1.5=23.3\ (m/s)" . Answer